OpenStudy (anonymous):

What is the exact value of the square root of the quantity 75 over 21. ? Simplify if possible.

OpenStudy (callisto):

\[\sqrt{\frac{75}{21}} = \sqrt{\frac{25\times 3}{7\times 3}}=\sqrt{\frac{25}{7}}= \sqrt{\frac{5^2}{7}}=?\]

OpenStudy (anonymous):

These are my choices 15 the square root of 3 all over 3. 15 the square root of 7 all over 7. 5 the square root of 7 all over 7. 5 the square root of 3 all over 3.

OpenStudy (callisto):

Hmm.. why don't you try first?

OpenStudy (anonymous):

I did... I got the same thing as you did but I don't see which answer that matches.

OpenStudy (anonymous):

Well, I got 5 over sqrt(7)

OpenStudy (callisto):

\[\frac{5}{\sqrt{7}} = \frac{5\sqrt{7}}{(\sqrt{7})^2}=?\]

OpenStudy (anonymous):

I don't understand

OpenStudy (callisto):

Multiply \[\sqrt7\] in both numerator and denominator so as to rationalise the expression. Now you need to simplify it, what do you get?